### v1

@@ 5.4 Model limitations
Hallucinations. CroissantLLM can hallucinate (Ji et al., 2023; Guerreiro et al., 2023a) and output factually incorrect data [^58] As an example, prompted with ”Which French club won the UEFA Champions League ?”, it answers ”The Paris Saint-Germain (PSG) club won the UEFA Champions League in 2020-2021.”, especially regarding complex topics. This is to be expected given the small model size, and hallucination rates seem inferior to most models of the same size category although no quantitative assessments have been conducted outside of MT-Bench experiments.
@@ 6 Foundation Model Transparency Index
In our efforts of transparency, we provide a summary of contributions for all authors and persons associated to the paper. Manuel, Patrick, Nuno, and Pierre belong to the core team and have participated throughout all steps of the project, from decision-making to report writing. Manuel coordinated the project and led the data collection, design decisions, and model evaluation efforts, and strongly participated in the model training. Pierre is the senior author of the project and was instrumental through constant feedback, project coordination, securing the compute grant, and design decisions. Patrick led the scaling law efforts and spearheaded the model training on distributed compute clusters. Nuno led the Chat and translation finetuning efforts, including constructing model finetuning pipelines and datasets, and gave constant feedback throughout the project. Pedro provided help on the development of the pre-training codebase and gave feedback on the pre-training stream of the work. João and Ricardo constructed the parallel data used for pre-training, which included efforts in both large-scale data collection and filtering. Duarte assisted with the fine-tuning efforts. António worked on base model finetuning on specific tasks and was in charge of the Chat model evaluation and the inference speed benchmark. Caio assisted with data collection efforts, and provided high-quality, extensive feedback and notes on the report. Nicolas assisted with data collection efforts and data scrapping. Antoni adapted the model to swiftly run on mobile devices. Gautier, Céline, François, André are senior researchers who provided valuable feedback and important guidance throughout the project and were instrumental in obtaining compute grants.
@@
This work is a joint collaboration of academic and industrial partners. On the academic side, core authors are affiliated with CentraleSupélec (Université Paris Saclay) and Instituto Superior Técnico de Lisboa, and other contributors are linked to Sorbonne Université and Imperial College London. On the industrial side, core authors receive funding from respectively Illuin Technology (Paris), Unbabel (Lisboa), Equall (New York, Lisboa, Paris). Training compute is obtained on the Jean Zay supercomputer operated by GENCI IDRIS through compute grant 2023-AD011014668R1 as well as on Adastra through compute grant AD010614770. Part of the work was also supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), DECOLLAGE (ERC-2022-CoG 101088763), Center for Responsible AI (2022-C05i0102-02), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020. Evaluation and finetuning experiments are done on the Ruche platform (Université Paris Saclay), as well as on privately owned compute centers.
@@
Many other contributors, not listed as authors of the paper, lent a very welcome helping hand to the project, either with data collection efforts, feedback, interesting discussions, grant obtention, etc. In no particular order, we extend our thanks to Hélène Rousset, Bruno Hays, Bilel Omrani, Sacha Muller, Elena Hinnekens and Robert Vesoul (Illuin Technology), Paul-Henry Cournède, Renaud Monnet, Géraud Faye (CentraleSupélec), Pierre-Etienne Devineau (DINUM), Louise-Anne Charles (BnF Datalab), the Unbabel TowerLLM team… Finally, we would like to warmly thank Stéphane Requena for allowing us access to Jeanzay, as well as Rémi Lacroix, and Etienne Malaboeuf for the debugging and technical support.
@@ D.1 Methodology
For generative tasks, we simply generate with the default settings, which is greedy sampling. We acknowlege CroissantLLM works best with higher temperature values but did not want to introduce stochasticity to the evaluation. We also limit each benchmark task to 2000 samples at most, to shorten evaluation time. All evaluations are reproducible through the code at https://github.com/ManuelFay/lm-evaluation-harness which will be merged in the main repository.
D.2 Perplexity results
Communication with the Pagnol model authors (Launay et al., 2021), enabled us to compute CroissantLLM perplexity on the same dataset and in the same conditions as the PagnolXL results shared in the paper. PagnolXL is a 1.5B decoder model, and is reported to obtain a French Treebank [^63] http://www.llf.cnrs.fr/en/Gens/Abeille/French-Treebank-fr.php in its SPMRL instance (Seddah et al., 2013) Perplexity value of 16.40 (test set), computed  [^64] https://huggingface.co/docs/transformers/perplexity with a rolling stride of 512 and a max sequence length of 2048 tokens. CroissantLLM obtains a perplexity value of 10.37, largely outperforming the biggest model in the Pagnol family.
D.3 MT-Bench
Turn 1 (Figure 11) and Turn 2 (Figure 12) results are shown. We notice small models, struggle with reasoning based tasks and contrained generation imposed by Turn 2 prompts.
@@ D.3 MT-Bench
D.4 Bias Assessment

### v4

@@ 5.4 Model limitations
Hallucinations. CroissantLLM can hallucinate (Ji et al., 2023; Guerreiro et al., 2023) and output factually incorrect data, [^60] As an example, prompted with ”Which French club won the UEFA Champions League ?”, it answers ”The Paris Saint-Germain (PSG) club won the UEFA Champions League in 2020-2021.” especially regarding complex topics. This is to be expected given the small model size, and hallucination rates seem inferior to most models of the same size category although no quantitative assessments have been conducted outside of MT-Bench experiments.
@@ 6 Foundation Model Transparency Index
In our efforts of transparency, we provide a summary of contributions for all authors and persons associated to the paper. Manuel, Patrick, Nuno, and Pierre belong to the core team and have participated throughout all steps of the project, from decision-making to report writing. Manuel coordinated the project and led the data collection, design decisions, and model evaluation efforts, and strongly participated in the model training. Pierre is the senior author of the project and was instrumental through constant feedback, project coordination, securing the compute grant, and design decisions. Patrick led the scaling law efforts and spearheaded the model training on distributed compute clusters. Nuno led the Chat and translation finetuning efforts, including constructing model finetuning pipelines and datasets, and gave constant feedback throughout the project. Pedro provided help on the development of the pre-training codebase and gave feedback on the pre-training stream of the work. João and Ricardo constructed the parallel data used for pre-training, which included efforts in both large-scale data collection and filtering. Duarte assisted with the fine-tuning efforts. António worked on base model finetuning on specific tasks and was in charge of the Chat model evaluation and the inference speed benchmark. Caio assisted with data collection efforts and provided high-quality, extensive feedback and notes on the report. Nicolas assisted with data collection efforts and data scrapping. Antoni adapted the model to swiftly run on mobile devices. Gautier, Céline, François, André are senior researchers who provided valuable feedback and important guidance throughout the project and were instrumental in obtaining compute grants.
@@
This work is a collaboration of academic and industrial partners. On the academic side, core authors are affiliated with CentraleSupélec (Université Paris Saclay) and Instituto Superior Técnico de Lisboa, and other contributors are linked to Sorbonne Université and Imperial College London. On the industrial side, core authors receive funding from respectively Illuin Technology (Paris), Unbabel (Lisboa), Equall (New York, Lisboa, Paris). Training compute is obtained on the Jean Zay supercomputer operated by GENCI IDRIS through compute grant 2023-AD011014668R1 as well as on Adastra through compute grant AD010614770. Part of the work was also supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), DECOLLAGE (ERC-2022-CoG 101088763), Center for Responsible AI (2022-C05i0102-02), and by Fundação para a Ciência e Tecnologia through contract UIDB/50008/2020. Evaluation and finetuning experiments are done on the Ruche platform (Université Paris Saclay), as well as on privately owned compute centers.
@@
Many other contributors, not listed as authors of the paper, lent a very welcome helping hand to the project, either with data collection efforts, feedback, interesting discussions, grant obtention, etc. In no particular order, we extend our thanks to Hélène Rousset, Bruno Hays, Bilel Omrani, Sacha Muller, Elena Hinnekens and Robert Vesoul (Illuin Technology), Paul-Henry Cournède, Renaud Monnet, Géraud Faye (CentraleSupélec), Pierre-Etienne Devineau (DINUM), Louise-Anne Charles (BnF Datalab), the Unbabel TowerLLM team. Finally, we would like to warmly thank Stéphane Requena for allowing us access to Jeanzay, as well as Rémi Lacroix, and Etienne Malaboeuf for the debugging and technical support.
@@ D.1 Methodology
For generative tasks, we simply generate with the default settings, which is greedy sampling. We acknowlege CroissantLLM works best with higher temperature values but did not want to introduce stochasticity to the evaluation. We also limit each benchmark task to 2000 samples at most, to shorten evaluation time. All evaluations are reproducible through the code at https://github.com/EleutherAI/lm-evaluation-harness.
D.2 MT-Bench
Turn 1 (Figure 11) and Turn 2 (Figure 12) results are shown. We notice small models, struggle with reasoning based tasks and contrained generation imposed by Turn 2 prompts. Figures 14 and 15 compare our results on small language models to other common bigger models. [^65] Results in French for models with sizes over 7B parameters were extracted from https://huggingface.co/datasets/bofenghuang/mt-bench-french and results in English are from https://huggingface.co/spaces/lmsys/mt-bench/tree/main/data/mt_bench/model_judgment
@@ D.2 MT-Bench
Models Wri Ro Reas Math Cod Ext STEM Hum Avg
CroissantLLMChat 5.32 5.35 1.9 1.16 1.8 1.4 4.55 3.75 3.15
TinyLlamaChat 5 4.1 1.45 1 2.1 1.6 3.9 4.85 3
Bloom 1B7 Chat 4.62 3.25 1.2 1.5 1.45 1.5 2.85 2.22 2.32
CMArkea BloomZ 3B 2.65 2.85 1.85 1.15 1.2 2.3 3.65 2.7 2.29
Vigostral 7B Chat 7.7 7.85 4.85 3.65 4.65 7.75 7.35 9.2 6.62
Vigogne 2 7b Chat 5.35 6.25 2.75 2.2 2.47 3.4 6.05 6.68 4.39
OpenHermes Mistral 7B 8.8 7.5 5.1 4.05 5.55 6.2 8.35 9.4 6.87
Vigogne 2 70B Chat 9.4 8.25 4.75 4.3 5.35 7.25 9.1 9.43 7.23
Mixtral 8x7b Instruct 9.65 8.88 6.95 4.95 4.6 8.55 9.5 9.6 7.84
Mistral Medium 9.6 9.05 5.4 6.1 7.35 9.25 9.3 9.75 8.23
GPT 3.5 Turbo 8.75 8.93 5.05 5.65 7.85 9.05 9.05 9.68 8
GPT 4 9.6 9.65 8.55 8.5 8.35 9.2 9.85 9.88 9.2
Table 14: French MT Bench Results Average of turn 1 and 2 of many supervised finetuned models
Models Wri Ro Reas Math Cod Ext STEM Hum Avg
CroissantLLMChat 5.1 3.6 2.4 1.1 1.8 1.3 5 5.85 3.27
TinyLlamaChat 6.25 5 3.3 1.35 2.1 1.5 4.82 6.3 3.83
Bloom 1B7 Chat 3.95 4.55 1.95 1.4 1.45 1.35 3.5 3.45 2.7
CMArkea BloomZ 3B 3.15 3.3 2.2 1.1 1.25 1.35 3.15 1.9 2.17
Llama 2 7B Chat 8.9 7.7 4.25 2.4 3 6.5 8.65 8.75 6.27
Vicuna 7B v1.3 8.1 7.45 4.65 2.3 3.55 5 7.82 9.1 6
Llama 2 13B Chat 8.85 7.5 5.1 3.45 3 6.92 8.62 9.75 6.65
Vicuna 13B v1.3 9.25 7.18 5.85 2.6 3.25 5.55 7.98 9.45 6.39
Vicuna 33B v1.3 9.5 8.45 6.65 3.15 3.35 7.1 8.98 9.8 7.12
Llama 2 70B Chat 9.3 7.5 5.8 3.3 3.15 7.25 8.93 9.62 6.86
GPT 3.5 Turbo 9.2 8.4 5.65 6.3 6.9 8.85 8.7 9.55 7.94
GPT 4 9.65 8.9 9 6.8 8.55 9.38 9.7 9.95 8.99
Table 15: English MT Bench Results Average of turn 1 and 2 of many supervised finetuned models
D.3 Bias Assessment